# Gravitational waves from dark domain walls

### v1

@@
Abstract
For most of cosmic history, the evolution of our Universe has been governed by the physics of a ‘dark sector’1, consisting of dark matter and dark energy, whose properties are only understood in a schematic way. The influence of these constituents is mediated exclusively by the force of gravity, meaning that insight into their nature must be gleaned from gravitational phenomena. The advent of gravitational-wave astronomy2, 3, 4 has revolutionised the field of black hole astrophysics, and opens a new window of discovery for cosmological sources. Relevant examples include topological defects, such as domain walls or cosmic strings5, 6, which are remnants of a phase transition. Here we present the first simulations of cosmic structure formation in which the dynamics of the dark sector introduces domain walls as a source of stochastic gravitational waves in the late Universe7, 8. We study in detail how the spectrum of gravitational waves is affected by the properties of the model, and extrapolate the results to scales relevant to the recent evidence for a stochastic gravitational wave background4. Our relativistic implementation9, 10, 11 of the field dynamics paves the way for optimal use of the next generation of gravitational experiments12, 13, 14, 15 to unravel the dark sector.
keywords:
@@ Abstract
The detection of the nanohertz stochastic gravitational wave background by the NANOGrav pulsar timing array4 opens a new window for studying gravitational systems at cosmological scales. The observed signal consists of correlated timing residuals of 68 millisecond pulsars that have now been measured over a 15-year time period. The amplitude of the timing residuals scales as $h_{c}^{2}f^{-3}$, where $f$ is the frequency and $h_{c}$ is the characteristic strain of the stochastic gravitational wave background16, i.e. the characteristic fractional change in the length of a hypothetical detector arm. An astrophysical background from supermassive black hole inspirals is predicted17 to yield $h_{c}\sim f^{-2/3}$ and hence a spectrum of timing residuals that tilts as $f^{-13/3}$. While supermassive black holes remain a viable explanation of the observed signal, a tilt between $f^{-2.6}$ and $f^{-3.8}$ is preferred by the data, which is in mild tension with the prediction. This has prompted a large number of investigations into alternative explanations, exploring possible phenomena in the dark sector or in the early Universe18, 19.
@@
One hypothetical source of gravitational waves are topological defects that are generically produced at phase transitions. Depending on the nature of the spontaneously broken symmetry, the defects can take the form of cosmic strings or domain walls20, 21. In particular, if the vacuum manifold splits into two (or more) isolated points, domain walls separate the regions that have settled into different vacua during the phase transition. The case where the phase transition happens in the early Universe at a high energy scale has been studied thoroughly in the recent literature22, 23, 24, 25, 19, owing to the dimensional argument that frequencies in the nanohertz range today correspond to the size of the causal horizon (the scale that sets the typical size of different vacuum domains) when the Universe had a temperature of several trillion degrees Kelvin.
@@
In this work instead, we consider domain walls appearing in the late-time Universe due to phase transitions in the dark energy field. This scenario has not yet been studied in the context of stochastic gravitational waves. We consider the (a)symmetron model26, a generalisation of the symmetron7, 8, a theory of gravity with an extra scalar degree of freedom. Such extra degrees of freedom are a general prediction of physics beyond the Standard Model. The (a)symmetron scalar field couples to the matter density and therefore mediates a fifth force that affects cosmological structure formation and is tightly constrained by local gravity experiments. When the local matter density falls below some critical threshold $\rho_{*}$, it triggers a $\mathbb{Z}_{2}$ phase transition, and the field settles into either of two minima. A crucial difference from the early Universe is that the late-time Universe has very large fluctuations in the local density, so that the phase transition happens in different places at different times. The resulting domains are separated by domain walls that are moulded by the sheets and filaments in the cosmic large-scale structure. This way, a large number of domain walls can be formed in a Hubble volume. Both their energy density and Compton scale are free parameters, but the former is often chosen well below the critical density such that the expansion rate of the Universe is not affected significantly. However, they can still produce a considerable stochastic gravitational wave signal if their time evolution is violent enough. In the case where the potential is not symmetric, the asymmetry tends to destabilise the domains that occupy the higher minimum. Spontaneous reconfiguration of domains then leads to the collision and annihilation of domain walls, adding a further source of gravitational waves that is qualitatively different from stable domain walls.
The asymmetron model depends on four parameters10: the Compton wavelength $L_{C}$, indicating the correlation length of the scalar field in vacuum; the cosmological redshift of the phase transition $z_{*}$, indicating the time when the background matter density is equal to the critical density $\rho_{*}$ of the phase transition; the average strength of the fifth force relative to gravity, $\bar{\beta}=(\beta_{+}+\beta_{-})/2$; and the respective difference $\Delta\beta=\beta_{+}-\beta_{-}$ in the two minima. The asymmetron reduces to the symmetron for the choice $\Delta\beta=0$, in which case the scalar potential is symmetric. We run two suites of simulations varying parameters about their respective values in a reference model, using Roman numerals \@slowromancapi@-\@slowromancapv@ for simulations in a cubic volume of $(742\,\mathrm{Mpc})^{3}$, and numerals i-v for the ones using a smaller volume of $(148\,\mathrm{Mpc})^{3}$. For quick reference, we use a subscript to indicate the parameter that was varied with respect to the reference model.
@@
The left panel of figure 1 shows the intensity of gravitational waves emanating from collapsing domain walls for the model \@slowromancapiii@${}_{\Delta\beta}$ that has an asymmetry of $10\%$ in the scalar field potential. In contrast, the right panel shows the gravitational wave pattern for a standard $\Lambda$-cold-dark-matter ($\Lambda$CDM) cosmological model without the scalar field. In this case, the main source of long-wavelength gravitational waves is the evolving large-scale structure.
Figure 2: Power spectra of the time derivative of the tensor perturbation, $\dot{h}_{ij}$, plotted as dimensionless energy density per logarithmic frequency interval. The conversion from Fourier wavenumber $k$ to frequency $f$ assumes the standard dispersion relation $k=2\pi f/c$. Results from the simulations in the larger volume, $(742\,\mathrm{Mpc})^{3}$, are shown in the left panel and the ones for the smaller volume, $(148\,\mathrm{Mpc})^{3}$, are shown in the right panel. The dotted line labelled “LSS” indicates the gravitational wave contribution from the evolution of large-scale structure as predicted from second-order perturbation theory. The short blue line labelled “SMBH”, inserted as a visual guide only, shows the expected slope for an astrophysical signal produced by supermassive black hole inspirals. This signal would however appear at very different frequencies and amplitudes.
Figure 2 shows the present-day ($z=0$) power spectra of gravitational waves, presented in units of dimensionless energy density per logarithmic frequency interval. There is a clear generation of power at the scale corresponding to the Compton wavelength of the theory with a power-law tail extending to lower frequencies. At extremely low frequency, below about $10^{-16}\,\mathrm{Hz}$, cosmic large-scale structure (LSS) becomes the dominant source of gravitational waves in most cases. The corresponding amplitude can be predicted from second-order perturbation theory and is indicated as dotted line in the figure, showing good agreement with our simulations in the regime where perturbation theory is valid. The somewhat larger amplitude we find for $\Lambda$CDM is due to higher-order nonlinearities and, at high wavenumber, discretisation noise in our numerical scheme. We checked that our results for the scalar field models are numerically converged.
The NANOGrav collaboration has presented confidence intervals for the spectral tilt by fitting a power law $\sim Af^{-\gamma}$ to the timing residuals4. We measure the spectral tilt $\gamma$ from our simulations, using the fact that the timing residuals scale as $\propto\Omega_{\mathrm{gw}}f^{-5}$. Our results are shown in figure 3. We note that all simulations display a plateauing spectral tilt that persists until the Compton scale. Furthermore, there is no strong indication for a dependence on $\bar{\beta}$ and $z_{*}$. Stable domain walls seem to generate slightly bluer spectra (lower $\gamma$) compared to the NANOGrav measurement, whereas collapsing and annihilating domain walls (models with $\Delta\beta>0$) push the tilt all the way to the red side of the NANOGrav posterior.
Figure 3: The spectral index $\gamma$ inferred from the power spectra shown in figure 2. The blue bands show the 68% and 95% confidence intervals reported for the NANOGrav observations that were however taken at much higher frequencies. The dotted horizontal line labelled SMBH indicates the expected spectral index for supermassive black hole inspirals, which would also appear at much higher frequencies.
To connect our results with the NANOGrav measurements, we need to extrapolate eight orders of magnitude towards smaller scales and about ten orders of magnitude in energy density $\Omega_{\mathrm{gw}}$. Such a huge extrapolation in parameter space comes with several caveats that we address partially, and they should be investigated further in the future. To get some analytic insight, it is useful to consider that the surface tension $\sigma$ of the domain walls scales as $\bar{\beta}^{2}(1+z_{*})^{6}L_{C}^{3}$ in vacuum. This sets the strength of the gravitational wave source and we expect that $\Omega_{\mathrm{gw}}\propto\sigma^{2}$. Our simulations indeed confirm this scaling with $\bar{\beta}$, while we find a slightly stronger-than-expected dependence $\Omega_{\mathrm{gw}}\sim(1+z_{*})^{14}$. The dependence on $L_{C}$ seems much more complicated than the naive consideration of $\sigma$ suggests. This may be due to the fact that domain walls tend to be aligned with sheets and filaments of dark matter, affecting the effective potential and hence the surface tension in a non-trivial way. We point out that the Compton scale $L_{C}$ sets the typical length scale over which the field can respond to local variations in the matter density.
Methods
@@
where $a$ denotes the scale factor, $c$ is the speed of light, $d\tau=a^{-1}dt$ is a conformal time element, $x^{i}$ are comoving Cartesian coordinates on the spatial hypersurface, and $h_{ij}$ is the transverse and traceless tensor perturbation containing the gravitational waves. A sum is implied over repeated indices. Our simulations also track scalar and vector perturbations of the metric11, but these are relevant mainly for matter dynamics and shall be omitted in the following discussion for brevity. In Fourier space, the tensor perturbations obey a damped oscillator equation that is driven by the spin-2 projection of the stress-energy tensor27,
@@
It is common practice to express the spectrum in terms of the energy density of gravitational waves28,
@@
When viewed on scales larger than the thickness of the domain wall, its stress-energy is simply proportional to the surface tension
@@
where one integrates between the two asymptotic vacuum values ϕ±subscriptitalic-$\phi_{\pm}$ on the two sides of the wall. At vanishing matter density we have $\sigma=12\sqrt{2}\,\Omega_{\mathrm{m}}^{2}H_{0}^{4}\bar{\beta}^{2}(1+z_{*})^{6}L_{C}^{3}$.
We have run [^1] The code is publicly available and can be downloaded at https://github.com/oyvach/AsGRD. three high-resolution simulations ($1280^{3}$ grid points, $742$ Mpc box size), labelled \@slowromancapi@-\@slowromancapiii@, and two lower-resolution simulations ($512^{3}$ grid points, $742$ Mpc box size), \@slowromancapiv@ and \@slowromancapv@. An additional five simulations roughly match the higher resolution but use a smaller volume ($256^{3}$ grid points, $148$ Mpc box size), and are labelled i-v. Among the larger volume simulations, the reference model \@slowromancapi@ has parameters ($L_{C},z_{*},\bar{\beta},\Delta\beta)=(1480\,\mathrm{kpc},0.1,8,0$). Models \@slowromancapii@-\@slowromancapv@ vary one parameter at a time, using $L_{C}=1111\,\mathrm{kpc}$, $\Delta\beta/\bar{\beta}=10\%$, $z_{*}=0$ and $\bar{\beta}=12$, respectively. The varied parameter is indicated as subscript in our figure labels for easy identification. In the smaller volume simulations, our reference model i has ($L_{C},z_{*},\bar{\beta},\Delta\beta)=(743\,\mathrm{kpc},-0.2,90,0$), and models ii-v use $L_{C}=801\,\mathrm{kpc}$, $\Delta\beta/\bar{\beta}=30\%$, $z_{*}=-0.13$ and $\bar{\beta}=120$, respectively.

### v2

@@
11institutetext: Institute of Theoretical Astrophysics, University of Oslo, Sem Sælands vei 13, 0371 Oslo, Norway
22institutetext: Institut für Astrophysik, Universität Zürich, Winterthurerstrasse 190, 8057 Zürich, Switzerland
33institutetext: CEICO, Institute of Physics of the Czech Academy of Sciences, Na Slovance 1999/2, 182 00, Prague 8, Czechia
Gravitational waves from dark domain walls
Ø. Christiansen oyvinch@fzu.cz1133    J. Adamek julian.adamek@uzh.ch22    F. Hassani farbod.hassani@astro.uio.no11    D. Mota , d.f.mota@astro.uio.no11
For most of cosmic history, the evolution of our Universe has been governed by the physics of a ‘dark sector’, consisting of dark matter and dark energy, whose properties are only understood in a schematic way. The influence of these constituents is mediated exclusively by the force of gravity, meaning that insight into their nature must be gleaned from gravitational phenomena. The advent of gravitational-wave astronomy has revolutionised the field of black hole astrophysics, and opens a new window of discovery for cosmological sources. Relevant examples include topological defects, such as domain walls or cosmic strings, which are remnants of a phase transition. Here we present the first simulations of cosmic structure formation in which the dynamics of the dark sector introduces domain walls as a source of stochastic gravitational waves in the late Universe. We study in detail how the spectrum of gravitational waves is affected by the properties of the model, and extrapolate the results to scales relevant to the recent evidence for a stochastic gravitational wave background. Our relativistic implementation of the field dynamics paves the way for optimal use of the next generation of gravitational experiments to unravel the dark sector.
Key Words.:
@@ 1 Introduction
The detection of the nanohertz stochastic gravitational wave background by the pulsar timing arrays (Agazie et al. 2023; Antoniadis et al. 2023; Reardon et al. 2023; Xu et al. 2023) opens a new window for studying gravitational systems at cosmological scales. The signal observed by the NANOGrav array (Agazie et al. 2023) consists of correlated timing residuals of 68 millisecond pulsars that have now been measured over a 15-year time period. The amplitude of the timing residuals scales as $h_{c}^{2}f^{-3}$, where $f$ is the frequency and $h_{c}$ is the characteristic strain of the stochastic gravitational wave background (Maiorano et al. 2021), i.e. the characteristic fractional change in the length of a hypothetical detector arm. An astrophysical background from supermassive black hole inspirals is predicted (Phinney 2001) to yield $h_{c}\sim f^{-2/3}$ and hence a spectrum of timing residuals that tilts as $f^{-13/3}$. While supermassive black holes remain a viable explanation of the observed signal, a tilt between $f^{-2.6}$ and $f^{-3.8}$ is preferred by the data, which is in mild tension with the prediction. This has prompted a large number of investigations into alternative explanations, exploring possible phenomena in the dark sector or in the early Universe (Agazie et al. 2023; Babichev et al. 2023).
@@
One hypothetical source of gravitational waves are topological defects that are generically produced at certain types of phase transitions. Depending on the nature of the spontaneously broken symmetry, the defects can take the form of cosmic strings or domain walls (Kibble 1976; Saikawa 2017; Vilenkin 1985; Vachaspati 2007). In particular, if the vacuum manifold splits into two (or more) isolated points, domain walls separate the regions that have settled into different vacua during the phase transition. The case where the phase transition happens in the early Universe at a high energy scale has been studied thoroughly in the recent literature (Ramazanov et al. 2022; Blasi et al. 2023; Li 2023; Babichev et al. 2023; Kitajima et al. 2023a; Gouttenoire & Vitagliano 2023; Kitajima et al. 2023b), owing to the dimensional argument that frequencies in the nanohertz range today correspond to the size of the causal horizon (the scale that sets the typical size of different vacuum domains) when the Universe had a temperature of several trillion degrees Kelvin.
In this work instead, we consider domain walls appearing in the late-time Universe due to phase transitions in the dark energy field. This scenario has not yet been studied in the context of stochastic gravitational waves. We consider the (a)symmetron model (Perivolaropoulos & Skara 2022), a generalisation of the symmetron (Hinterbichler & Khoury 2010; Hinterbichler et al. 2011), a theory of gravity with an extra scalar degree of freedom. Such extra degrees of freedom are a general prediction of physics beyond the Standard Model (Clifton et al. 2012). The symmetron can be regarded as a generic scalar effective field theory (EFT), with a potential parameterized as a quartic-order Taylor polynomial around the field origin. The quartic self-coupling term is chosen to be positive so that the potential is bounded from below, while the quadratic term is chosen to be negative in order to have a Higgs-like phase transition. The $\mathbb{Z}_{2}$ symmetry protects against large quantum corrections (Hinterbichler & Khoury 2010). The symmetron has a universal coupling to the matter sector through a conformal factor, which adds a coupling term to the effective mass of the scalar field. The model can thus be seen as a typical EFT limit for $\mathbb{Z}_{2}$ symmetric Horndeski theories in the case where the field $\phi/M_{\rm{pl}}$ is small so that the Taylor expansion can be truncated. In addition, we have added an odd term to the potential that explicitly breaks the $\mathbb{Z}_{2}$ symmetry, so that we can study the case where the vacuum degeneracy is lifted and one of the minima is favoured.
@@
The universal coupling to the matter sector mediates a fifth force that affects cosmological structure formation and is tightly constrained by local gravity experiments (Will 2014; Bertotti et al. 2003; Esposito-Farese 2004; Burrage et al. 2023; Llinares & Brax 2019). When the local matter density falls below some critical threshold $\rho_{*}$, it triggers the $\mathbb{Z}_{2}$ phase transition, and the field settles into either of two minima. A crucial difference from the early Universe is that the late-time Universe has very large fluctuations in the local density, so that the phase transition happens in different places at different times. The resulting domains are separated by domain walls that are moulded by the sheets and filaments in the cosmic large-scale structure. This way, a large number of domain walls can be formed in a Hubble volume. Both their energy density and Compton scale are free parameters, but the former is often chosen well below the critical density such that the expansion rate of the Universe is not affected significantly. However, they can still produce a considerable stochastic gravitational wave signal if their time evolution is violent enough. In the case where the potential is not symmetric, the asymmetry tends to destabilise the domains that occupy the higher minimum. Spontaneous reconfiguration of domains then leads to the collision and annihilation of domain walls, adding a further source of gravitational waves that is qualitatively different from stable domain walls.
In addition to being considered as a model for dark matter (Burrage et al. 2019; Käding 2023), the symmetron can also be motivated by the recent results of the Dark Energy Spectroscopic Instrument (DESI), which hint at a dynamical source for late-time cosmic acceleration (DESI Collaboration et al. 2024). In recent work (christiansen_environmental_2024-1) it was shown that a consistent evolution of the cosmic acceleration can be found using the symmetron together with a degravitation mechanism. Alternatively, one can still imagine obtaining a consistent accelerated expansion from the combination of the slow-rolling scalar with decreasing potential energy (and equation of state parameter $w\sim-1$) and the increasingly dominant domain walls ($w\sim-2/3$). The symmetron has also been linked to the detection of an anomalous Integrated Sachs-Wolfe (ISW) signal (Christiansen et al. 2024; Kovács et al. 2019), and to claims of large structure being in tension with the cosmological principle (Secrest et al. 2021, 2022; Peebles 2023; Lopez et al. 2024; Nadathur 2013).
@@
2 Gravitational wave spectra
@@
where $a$ denotes the scale factor, $c$ is the speed of light, $d\tau=a^{-1}dt$ is a conformal time element, $x^{i}$ are comoving Cartesian coordinates on the spatial hypersurface, and $h_{ij}$ is the transverse and traceless tensor perturbation containing the gravitational waves. A sum is implied over repeated indices. Our simulations also track scalar and vector perturbations of the metric (Adamek et al. 2016a), but these are relevant mainly for matter dynamics and shall be omitted in the following discussion for brevity. In Fourier space, the tensor perturbations obey a damped oscillator equation that is driven by the spin-2 projection of the stress-energy tensor (Adamek et al. 2016b),
@@
It is common practice to express the spectrum in terms of the energy density of gravitational waves (Allen & Romano 1999),
@@
3 Numerical simulations
The asymmetron model depends on four parameters (Christiansen et al. 2023): the Compton wavelength $L_{C}$, indicating the correlation length of the scalar field in vacuum; the cosmological redshift of the phase transition $z_{*}$, indicating the time when the background matter density is equal to the critical density $\rho_{*}$ of the phase transition; the average strength of the fifth force relative to gravity, $\bar{\beta}=(\beta_{+}+\beta_{-})/2$; and the respective difference $\Delta\beta=\beta_{+}-\beta_{-}$ in the two minima. The asymmetron reduces to the symmetron for the choice $\Delta\beta=0$, in which case the scalar potential is symmetric.
We run two suites of simulations [^1] The code (Christiansen et al. 2024, 2023; Adamek et al. 2016b; Daverio et al. 2016; Llinares & Mota 2013, 2014) is publicly available and can be downloaded at https://github.com/oyvach/AsGRD. varying parameters about their respective values in a reference model. Within the first suite, three high-resolution simulations ($1280^{3}$ grid points, $742$ Mpc box size) are labelled \@slowromancapi@-\@slowromancapiii@, and two lower-resolution simulations ($512^{3}$ grid points, $742$ Mpc box size) are labelled \@slowromancapiv@ and \@slowromancapv@. The five simulations within the second suite roughly match the higher resolution but use a smaller volume ($256^{3}$ grid points, $148$ Mpc box size), and are labelled i-v. Among the larger volume simulations, the reference model \@slowromancapi@ has parameters ($L_{C},z_{*},\bar{\beta},\Delta\beta)=(1480\,\mathrm{kpc},0.1,8,0$). Models \@slowromancapii@-\@slowromancapv@ vary one parameter at a time, using $L_{C}=1111\,\mathrm{kpc}$, $\Delta\beta/\bar{\beta}=10\%$, $z_{*}=0$ and $\bar{\beta}=12$, respectively. For quick reference, we use a subscript to indicate the parameter that was varied with respect to the reference model. In the smaller volume simulations, our reference model i has ($L_{C},z_{*},\bar{\beta},\Delta\beta)=(743\,\mathrm{kpc},-0.2,90,0$), and models ii-v use $L_{C}=801\,\mathrm{kpc}$, $\Delta\beta/\bar{\beta}=30\%$, $z_{*}=-0.13$ and $\bar{\beta}=120$, respectively. The choices of Compton wavelength are close to the resolution limit of these simulations, and we comment more on numerical resolution and convergence in appendix A. Smaller box sizes, allowing correspondingly smaller Compton wavelengths, become increasingly expensive to simulate due to the shorter dynamical timescales involved. The values of $\beta$ chosen are the largest that satisfy the code’s linearisation scheme (Christiansen et al. 2023). The small choices of $z_{*}\lesssim 0$ are largely unconstrained (Burrage et al. 2023).
The left panel of figure 1 shows the intensity of gravitational waves emanating from collapsing domain walls for the model \@slowromancapiii@Δβ that has an asymmetry of $10\%$ in the scalar field potential. In contrast, the right panel shows the gravitational wave pattern for a standard $\Lambda$-cold-dark-matter ($\Lambda$CDM) cosmological model without the scalar field. In this case, the main source of long-wavelength gravitational waves is the evolving large-scale structure.
Figure 2: Power spectra of the time derivative of the tensor perturbation, $\dot{h}_{ij}$, plotted as dimensionless energy density per logarithmic frequency interval. The conversion from Fourier wavenumber $k$ to frequency $f$ assumes the standard dispersion relation $k=2\pi f/c$. Results from the simulations in the larger volume, $(742\,\mathrm{Mpc})^{3}$, are shown in the left panel and the ones for the smaller volume, $(148\,\mathrm{Mpc})^{3}$, are shown in the right panel. The error bars indicate the 95% confidence intervals estimated from statistical fluctuations measured for each realisation within a narrow frequency band. The dotted line labelled “LSS” shows the gravitational wave contribution from the evolution of large-scale structure as predicted by second-order perturbation theory. The short blue line labelled “SMBH”, inserted as a visual guide only, indicates the expected slope for an astrophysical signal produced by supermassive black hole inspirals. This signal would however appear at very different frequencies and amplitudes.
Figure 2 shows the present-day ($z=0$) power spectra of gravitational waves, presented in units of dimensionless energy density per logarithmic frequency interval. There is a clear generation of power at the scale corresponding to the Compton wavelength of the theory with a power-law tail extending to lower frequencies. At extremely low frequency, below about $10^{-16}\,\mathrm{Hz}$, cosmic large-scale structure (LSS) becomes the dominant source of gravitational waves in most cases. The corresponding amplitude can be predicted from second-order perturbation theory and is indicated as dotted line in the figure, showing good agreement with our simulations in the regime where perturbation theory is valid. The somewhat larger amplitude we find for $\Lambda$CDM is due to higher-order nonlinearities and, at the highest wavenumbers, discretisation noise in our numerical scheme.
Figure 3: The spectral index $\gamma$ inferred from the power spectra shown in figure 2, with error bars indicating 95% confidence intervals. The blue bands show the 68% and 95% confidence intervals reported for the NANOGrav observations that were however taken at much higher frequencies. The dotted horizontal line labelled SMBH indicates the expected spectral index for supermassive black hole inspirals, which would also appear at much higher frequencies.
The NANOGrav collaboration has presented confidence intervals for the spectral tilt by fitting a power law $\sim Af^{-\gamma}$ to the timing residuals (Agazie et al. 2023). We measure the spectral tilt $\gamma$ from our simulations, using the fact that the timing residuals scale as $\propto\Omega_{\mathrm{gw}}f^{-5}$. Our results are shown in figure 3. We note that all simulations display a plateauing spectral tilt that persists up to the Compton scale. Furthermore, there is no strong indication for a dependence on $\bar{\beta}$ and $z_{*}$. Stable domain walls seem to generate slightly bluer spectra (lower $\gamma$) compared to the NANOGrav measurement, whereas collapsing and annihilating domain walls (models with $\Delta\beta>0$) push the tilt all the way to the red side of the NANOGrav posterior.
4 Discussion
To connect our results with the NANOGrav measurements, we need to extrapolate eight orders of magnitude towards smaller scales and about ten orders of magnitude in energy density $\Omega_{\mathrm{gw}}$. Such a huge extrapolation in parameter space comes with several caveats that we address partially, and they should be investigated further in the future. To get some analytic insight, it is useful to consider that, when viewed on scales larger than the thickness of the domain wall, its stress-energy is simply proportional to the surface tension
@@ 4 Discussion
where one integrates between the two asymptotic vacuum values ϕ±subscriptitalic-$\phi_{\pm}$ on the two sides of the wall. At vanishing matter density we have $\sigma=12\sqrt{2}\,\Omega_{\mathrm{m}}^{2}H_{0}^{4}\bar{\beta}^{2}(1+z_{*})^{6}L_{C}^{3}$. This sets the strength of the gravitational wave source and we expect that $\Omega_{\mathrm{gw}}\propto\sigma^{2}$. Our simulations indeed confirm this scaling with $\bar{\beta}$, while we find a slightly stronger-than-expected dependence $\Omega_{\mathrm{gw}}\sim(1+z_{*})^{14}$. The dependence on $L_{C}$ seems much more complicated than the naive consideration of $\sigma$ suggests. This may be due to the fact that domain walls tend to be aligned with sheets and filaments of dark matter, affecting the effective potential and hence the surface tension in a non-trivial way. We point out that the Compton scale $L_{C}$ sets the typical length scale over which the field can respond to local variations in the matter density.